# Lindbladian dynamics with loss of quantum jumps

### v2

@@ Acknowledgements.
$\rho=\left(\begin{matrix}\rho_{ee}&\rho_{eg}\\ \rho_{ge}&\rho_{gg}\end{matrix}\right)\ \ \to\ \ |\rho\rangle=(\rho_{ee},\ \rho_{eg},\ \rho_{ge},\ \rho_{gg})^{T}.$ (S10)
@@
$\frac{d}{dt}\,|\rho\rangle\,=\,\mathrm{L}_{\eta}\,|\rho\rangle,$ (S11)
@@
$\begin{split}\mathrm{L}_{\eta}=&-i\mathrm{H}\otimes\mathrm{I}+i\mathrm{I}\otimes\mathrm{H}^{\mathrm{T}}-\frac{\gamma}{2}(\mathrm{L}^{\dagger}\mathrm{L}\otimes\mathrm{I}+\mathrm{I}\otimes\mathrm{L}^{\mathrm{T}}\mathrm{L}^{*})\\ &+(1-\eta)\gamma\mathrm{L}\otimes\mathrm{L}^{*}+\eta\gamma<\mathrm{L}^{\dagger}\mathrm{L}>,\\ \end{split}$ (S12)
@@
$\mathrm{H}=\left(\begin{matrix}0&J\\ J&0\end{matrix}\right),\ \ \mathrm{L}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\ \ \mathrm{I}=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right).$ (S13)
@@
$\frac{d}{dt}\left(\begin{matrix}\rho_{ee}\\ \rho_{eg}\\ \rho_{ge}\\ \rho_{gg}\end{matrix}\right)=\left(\begin{matrix}-iJ(\rho_{ge}-\rho_{eg})+\eta\gamma\rho_{ee}^{2}-\gamma\rho_{ee}\\ -iJ(\rho_{gg}-\rho_{ee})+\eta\gamma\rho_{eg}\rho_{ee}-0.5\gamma\rho_{eg}\\ iJ(\rho_{gg}-\rho_{ee})+\eta\gamma\rho_{ge}\rho_{ee}-0.5\gamma\rho_{ge}\\ iJ(\rho_{ge}-\rho_{eg})+\eta\gamma\rho_{gg}\rho_{ee}+(1-\eta)\gamma\rho_{ee}.\end{matrix}\right)$ (S14)
@@
S3. Monte Carlo wave-function method for NLME.
@@
S4. Numerically confirm the validity of NLME
@@
Figure S1: The simulation of excited state probability from three methods. The system is described by Eq (9) in the main text with $\gamma=0.5$ and $\eta=0.2$. The atom is initialized in the excited state $|e\rangle$. The excited state probability is simulated by the methods “QT1”(blue solid line), “QT2”(green dashed line) and “NLME”(red cross).
(1) According to the proposal of experimental implementation as shown in Fig. 1(a) in the main text, the postselection processes need to monitor dissipations and collect the experimental data without detectors clicking. The most faithful simulation is based on the original quantum trajectory method with postselecting the trajectories, which we mark as the method “QT1”. The process is as follows. First, We simulate the system evolution by Eq. (S1). For every decay channel in every time step $\delta t$, for example, $L_{\mu}$ in time $t\sim t+\delta t$, we use a random number $r_{1}\in[0,1)$ to choose the evolution processes. If $r_{1}<\delta p_{\mu}$, the quantum jump occurs and evolution is $|\phi(t+\delta t)\rangle=\mathcal{N}L_{\mu}|\phi(t)\rangle$. If $r_{1}\geq\delta p_{\mu}$, the non-unitary evolution occurs, i.e., $|\phi(t+\delta t)\rangle=\mathcal{N}exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)|\phi(t)\rangle$. Second, if quantum jumps occur, we use a second random number $r_{2}\in[0,1)$ to decide whether the jump will be detected. If $r_{2}<\eta_{\mu}$, jumps is detected and we discard this trajectory. Only the trajectories with no jumps being detected are saved. Third, we use the saved trajectories to calculate the observation average values.
(2) Discarding quantum jumps from postselection is equivalent to increasing (decreasing) the probability of the processes of non-unitary evolution (quantum jumps). The second simulation method is just adjusting the probability accroding to the postselection effciency $\eta_{\mu}$ in the original quantum trajectory method as shown in by Eq. (S2) and Eq. (S3). Therefore, we can calculate the postselection dynamics by the ensemble of trajectories produced by Eq. (S4) without further postselecting the trajectories. We mark the second method as “QT2”.
(3) We use the method ”QT2” to derive the nonlinear Lindblad master equation (NLME) in Eq. (S9). However, the NLME as the ordinary differential equation can be directly simulated by some numerical methods such as Runge-Kutta methods. We mark the simulation directly from the NLME as the method “NLME”.
We simulate the postselection dynamics of the two-level atom with monitoring its spontaneous emission in Eq (9) in the main text by the above three methods. The three curves of time evolution of excited state probability are almost the same as shown in Fig. S1. This confirms the validity of NLME.
S5. The Gaussian state method for free fermions.
@@
$\hat{d}^{\dagger}_{\bm{v}}=(c^{\dagger}_{1},c^{\dagger}_{2},\cdots c^{\dagger}_{L})\cdot\bm{v}=\sum_{i=1}^{L}v_{i}c^{\dagger}_{i},$ (S15a)
$\hat{d}_{\bm{v}}=\bm{v}^{\dagger}\cdot(c_{1},c_{2},\cdots c_{L})^{\mathrm{T}}=\sum_{i=1}^{L}v_{i}^{*}c_{i}.$ (S15b)
@@
$\mathcal{N}|V\rangle=\mathcal{N}\prod_{n=1}^{N}(\sum_{j=1}^{n}R_{jn}\hat{d}^{\dagger}_{\bm{q}_{j}})|0\rangle=\prod_{n=1}^{N}\hat{d}^{\dagger}_{\bm{q}_{j}}|0\rangle=|Q\rangle,$ (S16)
@@
$S_{L}(a,b)=-\sum_{i}\left[\lambda_{i}ln\lambda_{i}+(1-\lambda_{i})ln(1-\lambda_{i})\right],$ (S17)
@@
$\hat{J}_{M}=(c^{\dagger}_{1},c^{\dagger}_{2},\cdots c^{\dagger}_{L})M(c_{1},c_{2},\cdots c_{L})^{\mathrm{T}},$ (S18)
@@
$\begin{split}\mathcal{N}e^{\hat{J}_{M}}|V\rangle&=\mathcal{N}\prod_{n=1}^{N}\left(\sum_{j=1}^{L}e^{\hat{J}_{M}}c^{\dagger}_{j}e^{-\hat{J}_{M}}V_{jn}\right)|0\rangle\\ &=\mathcal{N}\prod_{n=1}^{N}\left(\sum_{m=1}^{L}c^{\dagger}_{m}(e^{M}V)_{mn}\right)|0\rangle=\mathcal{N}|e^{M}V\rangle.\\ \end{split}$ (S20)
@@
$\mathcal{N}\hat{d}^{\dagger}_{\bm{\alpha}}|V\rangle=\mathcal{N}|\bm{\alpha}\otimes V\rangle:=\mathcal{N}|\{\bm{\alpha},\bm{v}_{1},\bm{v}_{2}\cdots\bm{v}_{N}\}\rangle.$ (S21)
@@
$\hat{d}_{\bm{\alpha}}\hat{d}^{\dagger}_{\bm{v}}=(\sum_{i}\alpha^{*}_{i}c_{i})(\sum_{j}v_{j}c^{\dagger}_{j})=\bm{\alpha}^{\dagger}\cdot\bm{v}-\hat{d}^{\dagger}_{\bm{v}}\hat{d}_{\bm{\alpha}},$ (S22)
@@
$\hat{d}_{\bm{\alpha}}|V\rangle=\sum_{l=1}^{N}(-1)^{l-1}(\bm{\alpha}^{\dagger}\cdot\bm{v}_{l})\prod_{n\neq l}\hat{d}^{\dagger}_{\bm{v}_{n}}|0\rangle,$ (S23)
@@
$\mathcal{N}\hat{d}_{\bm{\alpha}}|V\rangle=\mathcal{N}\ |\ {\rm ker}\left((\bm{\alpha}\otimes{\rm ker}(V^{\dagger}))^{\dagger}\right)\ \rangle,$ (S24)
@@
S6. Distribution and scaling of entangled entropy
The postselected skin effect profoundly affects the distribution and scaling behaviors of TAEE by the tanh-fitting steady state, where the particle number only changes rapidly in the middle of the chain. Near the ends of a long chain, the local site is nearly fully occupied ($n_{x}\approx 1$) or near vacuum ($n_{x}\approx 0$), which contributes little to entangled entropy. This makes the entropy distribution $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ increase with a small $\Delta$ while saturate for large $\Delta$, as shown in Fig. S2(a).
Figure S2: Distribution and scaling of steady-state entangled entropy. (a) Steady-state entanglement entropy $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ as a function of the bipartition size $\Delta$ in a half-filled $150$-site chain with $\eta=0.6$ and $\gamma$ from 00 to $1$ in the log-log plot. (b) Half-chain entanglement entropy $\overline{S}_{L}(x_{C},x_{C}+L/2)$ scaling with the half-filled chain length $L$ in the log-log plot. The entropy has a two-segment scaling behavior: rapid growth in small $L$ and saturation in large $L$. With decreasing $\gamma$, the rapid growth segment becomes longer. Especially in $\gamma=0$, the growth segment increases to the whole chain, which implies a volume law of entropy in the non-dissipative free systems.
The steady state’s scaling behavior of the half-chain TAEE $\overline{S}_{L}(x_{C},x_{C}+L/2)$ also presents a two-segment structure. When increasing $L$, the population distribution of the whole chain becomes a larger part of the tanh function Eq. (22) in the main text, as shown in Fig. 2(b) in the main text, and meanwhile the end-site particle number $n_{L}$ decreases rapidly from $0.5$ and then approaches slowly to 00. This leads to $\overline{S}_{L}(x_{C},x_{C}+L/2)$ having a nearly algebraic growth in small $L$ and saturation in large $L$, as shown in Fig. S2(b).

### v3

@@ Appendix B Vectorization treatment of the two-level atom
$\rho=\left(\begin{matrix}\rho_{ee}&\rho_{eg}\\ \rho_{ge}&\rho_{gg}\end{matrix}\right)\ \ \to\ \ |\rho\rangle=(\rho_{ee},\ \rho_{eg},\ \rho_{ge},\ \rho_{gg})^{T}.$ (36)
@@
$\frac{d}{dt}\,|\rho\rangle\,=\,\mathrm{L}_{\eta}\,|\rho\rangle,$ (37)
@@
$\begin{split}\mathrm{L}_{\eta}=&-i\mathrm{H}\otimes\mathrm{I}+i\mathrm{I}\otimes\mathrm{H}^{\mathrm{T}}-\frac{\gamma}{2}(\mathrm{L}^{\dagger}\mathrm{L}\otimes\mathrm{I}+\mathrm{I}\otimes\mathrm{L}^{\mathrm{T}}\mathrm{L}^{*})\\ &+(1-\eta)\gamma\mathrm{L}\otimes\mathrm{L}^{*}+\eta\gamma<\mathrm{L}^{\dagger}\mathrm{L}>,\\ \end{split}$ (38)
@@
$\mathrm{H}=\left(\begin{matrix}0&J\\ J&0\end{matrix}\right),\ \ \mathrm{L}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\ \ \mathrm{I}=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right).$ (39)
@@
$\frac{d}{dt}\left(\begin{matrix}\rho_{ee}\\ \rho_{eg}\\ \rho_{ge}\\ \rho_{gg}\end{matrix}\right)=\left(\begin{matrix}-iJ(\rho_{ge}-\rho_{eg})+\eta\gamma\rho_{ee}^{2}-\gamma\rho_{ee}\\ -iJ(\rho_{gg}-\rho_{ee})+\eta\gamma\rho_{eg}\rho_{ee}-0.5\gamma\rho_{eg}\\ iJ(\rho_{gg}-\rho_{ee})+\eta\gamma\rho_{ge}\rho_{ee}-0.5\gamma\rho_{ge}\\ iJ(\rho_{ge}-\rho_{eg})+\eta\gamma\rho_{gg}\rho_{ee}+(1-\eta)\gamma\rho_{ee}.\end{matrix}\right)$ (40)
@@
Appendix C Monte Carlo wave-function method for NLME
@@
Appendix D Numerically confirm the validity of NLME
@@
Figure 5: The simulation of excited state probability from three methods. The system is described by Eq (9) in the main text with $\gamma=0.5$ and $\eta=0.2$. The atom is initialized in the excited state $|e\rangle$. The excited state probability is simulated by the methods “QT1”(red cross), “QT2”(green dashed line) and “NLME”(blue solid line).
(1) According to the proposal of experimental implementation as shown in Fig. 1(a) in the main text, the postselection processes need to monitor dissipations and collect the experimental data without detectors clicking. The most faithful simulation is based on the original quantum trajectory method with postselecting the trajectories, which we mark as the method “QT1”. The process is as follows. First, We simulate the system evolution by Eq. (27). For every decay channel in every time step $\delta t$, for example, $L_{\mu}$ in time $t\sim t+\delta t$, we use a random number $r_{1}\in[0,1)$ to choose the evolution processes. If $r_{1}<\delta p_{\mu}$, the quantum jump occurs and evolution is $|\phi(t+\delta t)\rangle=\mathcal{N}L_{\mu}|\phi(t)\rangle$. If $r_{1}\geq\delta p_{\mu}$, the non-unitary evolution occurs, i.e., $|\phi(t+\delta t)\rangle=\mathcal{N}exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)|\phi(t)\rangle$. Second, if quantum jumps occur, we use a second random number $r_{2}\in[0,1)$ to decide whether the jump will be detected. If $r_{2}<\eta_{\mu}$, jumps is detected and we discard this trajectory. Only the trajectories with no jumps being detected are saved. Third, we use the saved trajectories to calculate the observation average values.
(2) Discarding quantum jumps from postselection is equivalent to increasing (decreasing) the probability of the processes of non-unitary evolution (quantum jumps). The second simulation method is just adjusting the probability accroding to the postselection effciency $\eta_{\mu}$ in the original quantum trajectory method as shown in by Eq. (28) and Eq. (29). Therefore, we can calculate the postselection dynamics by the ensemble of trajectories produced by Eq. (30) without further postselecting the trajectories. We mark the second method as “QT2”.
(3) We use the method ”QT2” to derive the nonlinear Lindblad master equation (NLME) in Eq. (35). However, the NLME as the ordinary differential equation can be directly simulated by some numerical methods such as Runge-Kutta methods. We mark the simulation directly from the NLME as the method “NLME”.
We simulate the postselection dynamics of the two-level atom with monitoring its spontaneous emission in Eq (9) in the main text by the above three methods. The three curves of time evolution of excited state probability are almost the same as shown in Fig. 5. This confirms the validity of NLME.
Appendix E The Gaussian state method for free fermions
@@
$\hat{d}^{\dagger}_{\bm{v}}=(c^{\dagger}_{1},c^{\dagger}_{2},\cdots c^{\dagger}_{L})\cdot\bm{v}=\sum_{i=1}^{L}v_{i}c^{\dagger}_{i},$ (41a)
$\hat{d}_{\bm{v}}=\bm{v}^{\dagger}\cdot(c_{1},c_{2},\cdots c_{L})^{\mathrm{T}}=\sum_{i=1}^{L}v_{i}^{*}c_{i}.$ (41b)
@@
$\mathcal{N}|V\rangle=\mathcal{N}\prod_{n=1}^{N}(\sum_{j=1}^{n}R_{jn}\hat{d}^{\dagger}_{\bm{q}_{j}})|0\rangle=\prod_{n=1}^{N}\hat{d}^{\dagger}_{\bm{q}_{j}}|0\rangle=|Q\rangle,$ (42)
@@
$S_{L}(a,b)=-\sum_{i}\left[\lambda_{i}ln\lambda_{i}+(1-\lambda_{i})ln(1-\lambda_{i})\right],$ (43)
@@
$\hat{J}_{M}=(c^{\dagger}_{1},c^{\dagger}_{2},\cdots c^{\dagger}_{L})M(c_{1},c_{2},\cdots c_{L})^{\mathrm{T}},$ (44)
@@
$\begin{split}\mathcal{N}e^{\hat{J}_{M}}|V\rangle&=\mathcal{N}\prod_{n=1}^{N}\left(\sum_{j=1}^{L}e^{\hat{J}_{M}}c^{\dagger}_{j}e^{-\hat{J}_{M}}V_{jn}\right)|0\rangle\\ &=\mathcal{N}\prod_{n=1}^{N}\left(\sum_{m=1}^{L}c^{\dagger}_{m}(e^{M}V)_{mn}\right)|0\rangle=\mathcal{N}|e^{M}V\rangle.\\ \end{split}$ (46)
@@
$\mathcal{N}\hat{d}^{\dagger}_{\bm{\alpha}}|V\rangle=\mathcal{N}|\bm{\alpha}\otimes V\rangle:=\mathcal{N}|\{\bm{\alpha},\bm{v}_{1},\bm{v}_{2}\cdots\bm{v}_{N}\}\rangle.$ (47)
@@
$\hat{d}_{\bm{\alpha}}\hat{d}^{\dagger}_{\bm{v}}=(\sum_{i}\alpha^{*}_{i}c_{i})(\sum_{j}v_{j}c^{\dagger}_{j})=\bm{\alpha}^{\dagger}\cdot\bm{v}-\hat{d}^{\dagger}_{\bm{v}}\hat{d}_{\bm{\alpha}},$ (48)
@@
$\hat{d}_{\bm{\alpha}}|V\rangle=\sum_{l=1}^{N}(-1)^{l-1}(\bm{\alpha}^{\dagger}\cdot\bm{v}_{l})\prod_{n\neq l}\hat{d}^{\dagger}_{\bm{v}_{n}}|0\rangle,$ (49)
@@
$\mathcal{N}\hat{d}_{\bm{\alpha}}|V\rangle=\mathcal{N}\ |\ {\rm ker}\left((\bm{\alpha}\otimes{\rm ker}(V^{\dagger}))^{\dagger}\right)\ \rangle,$ (50)